OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 + 2(x − 1) = 5x − (x + 10) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (ajprincess):

R u familiar with distributive property?

OpenStudy (anonymous):

not really

OpenStudy (ajprincess):

a(b+c)=a*b+a*c =ab+ac This is it.

OpenStudy (ajprincess):

Nw can u expand 2(x − 1) using it?